OpenStudy (anonymous):

v v 1 -- - -- = -- solve v= 21 35 21

OpenStudy (anonymous):

LCM=3*5*7 (5v-3v)/(3*5*7)=1/21 2v=(3*5*7)/3*7 v=5/2